-x^2(x+2)(x^2-1) how can i find the x intercepts. I get lost factoring! Help
Answer by ewatrrr(24785)

 
Hi
Re: TY, graph was for demonstration only, in this case:
Note: The factor theorem states that a polynomial f(x) has a factor (x − k)if and only if f(k) = 0
IF -x^2(x+2)(x-1)(x+1)= 0 || IF f(k) = 0, then the resulting "k(s)" are x-intercepts!
-x^2 = 0 ||x = 0
0r (x+2) = 0 || x = -2
0r (x-1) = 0 || x = 1
0r (x+1) = 0 || x = -1
-x^2(x+2)(x^2-1) = -x^2(x+2)(x-1)(x+1) |||x-intercepts are -2,-1,0,1
Might recommend checking out downloading this FREE graphing software

www.padowan.dk